A spring of force constant K is cut into two parts of length in the ratio 1:2. spring constant of the larger one is

Grade:11

1 Answers

Arun
5 years ago
 
let length of the spring having spring constant k is L. now as spring is cut into 2 pieces .
                                          L1= L/3
                                           L2=2 x L/3   (as one piece is double of other in length)
 kL=k1 L1=k2 L2
    kL=k2(2L/3)
     3k/2=k2
 therefore k2=3k/2
 
Regards
Arun